OpenStudy (anonymous):

5x^2 =10 solve using radicals,rstionalize all denominators express complex numbers in terms of i or use a comma to separate answers if needed

OpenStudy (anonymous):

divide both sides by 5 x^2 = 2 now take square roots of both sides

OpenStudy (anonymous):

is that ok?

OpenStudy (anonymous):

im just not sure how to do this one it leaves the answer open to too many possibilities

OpenStudy (anonymous):

well the square root of x^2 is , by definition, x and there are 2 sq roots of 2:

OpenStudy (anonymous):

\[\sqrt{2} , -\sqrt{2}\]

OpenStudy (anonymous):

there are no complex numbers in this result - just two radicals

OpenStudy (anonymous):

i see makes sense

OpenStudy (anonymous):

if this were a grapg what would the x intercepts be